A uniform rod rotates in a horizontal plane about a vertical axis through a point


which is 2 m away from its center. The rod is 8 m long, weighs 8 N, and rotates at 100


rev/minute. Calculate


(a) Its rotational inertia about the axis of rotation.


(b) The magnitude of its angular momentum about the axis.

1
Expert's answer
2021-12-17T18:16:04-0500

(a) Its rotational inertia about the axis of rotation:


"I=\frac1{12}mL^2+md^2,\\\space\\
I=\frac1{12}·\frac8{9.8}·8^2+\frac8{9.8}·2^2=7.6\text{ kg}·\text{m}^2."

(b) The magnitude of its angular momentum about the axis:


"L=I\omega=I·2\pi n=4775\text{ kg}·\text{rad/s}."